# AlphaZero Neural Scaling and Zipf’s Law: a Tale of Board Games and Power Laws

### v1

@@ 6 Analyzing how loss scales with Zipf’s law
In agreement with the quantization model (Michaud et al., 2023), we find that AlphaZero models reduce their loss on game states in descending order of frequency. In Fig. 4 we plot the value-head loss on Connect Four training data in the limit of abundant compute, tracking the loss of each agent on its self-play games.
Figure 4: Connect Four value-loss [^3] Note that AlphaZero’s value head is not equivalent to the value net of actor-critic methods, see appendix D.scaling with rank. A: Average loss on each agent’s own training data. Loss steadily increases with rank. Larger agents achieve better loss. B: Loss on a dataset annotated by a game solver. Larger models are closer to the ground truth values of optimal play. The overall trend of loss increasing with rank is not trivial, considering that the complexity of states decreases with rank. C: The time needed to evaluate states using alpha-beta pruning (Pons, 2015) drops exponentially with rank (mean and standard deviation are geometric).
Comparison to LLMs
The quantization hypothesis assumes LLMs learn tasks in a binary way, sharply reducing the loss on each task in descending order of frequency. As a function of neural net size $N$, this assumed loss function is a step function that changes from low to high loss at rank $n\propto N$ (Michaud et al., 2023). In comparison, we find that AlphaZero loss increases smoothly with rank rather than resembling a step function, likely because board positions are not independent task quanta.
In Fig. 4B we plot the value loss $L_{\rm value}=\left(z-v\right)^{2}$ on ground-truth labels $z$, calculated by an alpha-beta pruning solver (Pons, 2015). We see that larger agents have significantly better ground-truth loss over smaller agents, confirming that increasing model size improves the absolute value loss. Ground-truth value loss is directly related to better performance, similar to test loss in LLMs that is known to correlate with accuracy on downstream tasks (Brown et al., 2020; Srivastava et al., 2022; Wei et al., 2022).
@@
It is surprising that absolute value loss (Fig. 4B) tends to increase with state rank, since higher rank states are easier to model, meaning loss increases as state complexity decreases. Low-loss states are the hardest to model using conventional methods, as we show in Fig. 4C for alpha-beta pruning (Knuth & Moore, 1975). The most frequent states have a larger state tree branching out from them, making it harder for search-based algorithms like MCTS to tackle them. As a result, the CPU time required to calculate the ground truth value of a state decreases steadily with rank on an exponential scale.
Figure 5: Inverse scaling in AlphaZero. A: Agents playing Checkers and Oware follow a size scaling law that abruptly changes direction, when large models fail to utilize their capacity. The scaling curve does not flip due to approaching the perfect-play ceiling: training a suite of Oware agents with different hyperparameters (light green), they extend to higher Elo scores. Error bars are one standard deviation. B: Oware games played by AlphaZero follow Zipf’s law, interrupted by a small plateau. This is caused by high-frequency late-game states, present due to the unusual tree-structure of Oware. Compare Fig. 6. C: Checkers shares the same tree structure, but the effect on the Zipf curve is less visible.
@@ 7 Connecting inverse scaling to the game structure
Inverse scaling, i.e. performance scaling negatively with training resources, is known to occur in LLM training (McKenzie et al., 2023). RL models can similarly fail to scale reliably, sometimes decreasing in performance as neural-network size and compute budget are increased. We see a clear example of inverse scaling for AlphaZero agents trained on Checkers and Oware, as shown in Fig. 5A. These agents follow a steady scaling curve that abruptly breaks down beyond a certain neural network size, where Elo starts to scale negatively with size.
@@ 7.1 Inverse scaling coincides with a frequency-distribution anomaly
Figure 6: Turn distribution anomaly coincides with inverse scaling. A: When the game tree spreads exponentially, as for Connect Four and Pentago, turn number correlates with state frequency. However, in games with inverse scaling (Oware and Checkers) the game tree converges to a comparatively small set of endings, resulting in high-frequency late-game states. B: The fraction of late-game states (above dashed line) rises either sharply or gradually with rank.
@@
States above and below the dashed line in Fig. 6A are clearly distinguishable when visualized, see examples in appendix E. Late-game states above the dashed line are positions close to the end of a game, when the board is mostly empty, while early-game states below the line are opening moves, i.e. small variations of the initial board state.
@@
The prevalence of high-frequency late-game states in Oware and Checkers data is a direct result of the game rules, which shape the game-tree structure. In both games, pieces are successively removed from the board, ending the game when one of the players has no pieces left (infinite-loop draws are rarer). This win condition produces game trees that expand and then contract: starting from a predetermined position, expanding exponentially every move, but then contracting again towards the end of the game as the number of pieces decreases and with it the number of possible board-state permutations. The probability that a game ends in a frequently-visited board state is high.
@@ 7.2 Relation to inverse scaling
As we show in Fig. 6, Late-game states (above the dashed line) dominate the Oware distribution at higher ranks, in a sharp transition from $0\%$ to $>80\%$ of all states. This transition is drastic enough to visibly skew the frequency distribution seen in Fig. 5B, bending the Zipf power law. The distribution shift causes Oware value loss to spike, as we show in Fig. 7A. The loss of all but the largest models jumps by up to an order of magnitude at the transition point, showing that models that fit early-game states well have a hard time fitting late-game states. This counter-intuitive behavior of getting worse loss on easier states, which we discussed in section 6, is even more surprising here: Oware states after the transition point are predominantly end-game configurations, and should be significantly easier to model than early-game states using search algorithms.
The quantization model provides an explanation for the sharp rise in loss. As models fit the most frequent states, smaller models do not have enough capacity to additionally model less frequent states accurately. The large qualitative difference between early- and late-game states makes the difference in loss especially pronounced. Interestingly, larger agents converge on a smooth loss curve similar to that of Connect Four, at the cost of worse loss on early-game states.
@@ 7.3 Why do large models overfit?
The degradation of large model performance on early-game states is an example where RL scaling diverges from LLM scaling theory. The quantization model applied to LLMs explains how larger models could fit more tasks, without forgetting higher-frequency tasks. In RL, we see that fitting less-frequent states coincides with lower performance on the most frequent states, when states split into two distinct groups. The difference between LLM and RL scaling might be caused by a shift of the target distribution during training, which is known to cause training failure, e.g. through dormant neurons (Sokar et al., 2023).
Figure 8: Large models forget knowledge on early-game states. A: The loss of the largest Oware agent improves (on average) during training. B: On early-game states, loss initially goes down, then stops improving and eventually worsens towards the end of training. C: Loss on late-game states steadily improves during training. The fact that large nets fit early-game states but then forget them suggests target non-stationarity might be the cause of inverse scaling.
@@
A possible explanation of this selective forgetting is a selective shift in the distribution of the target labels (the value of each state). Sokar et al. (2023) have shown that changes to the target distribution in RL can harm performance by creating dormant neurons. Unlike early-game states, late-game states have relatively fixed label distributions since the outcome is already decided by that point of the game. A hypothetical mechanism of inverse scaling then follows: 1) Models fit states by order of frequency, gradually expanding the fraction of late-game states in the pool of learned states. 2) As new states are optimized, higher-frequency states learned earlier can be forgotten. 3) In Oware, late-game states are immune to forgetting due to their strongly fixed labels. 4) Large Oware models are trapped in local optima, fitting many end-game states well but forgetting a smaller fraction of early-game states.
@@
We note a certain similarity between this inverse scaling phenomenon and one of the causes for inverse scaling identified by McKenzie et al. (2023), namely distractor tasks. Some LLM benchmarks require mastering a set of tasks $T$, but as model scale increases, models first learn a set of easier tasks $D$ that harm performance, leading to inverse scaling. Another similar concept is competing tasks, which can cause inverse scaling in a supervised learning setting (Ildiz et al., 2024). While the concept of competing/distractor tasks bears resemblance to the Oware case, the mechanism behind these phenomena is different. These examples of supervised learning inverse scaling are driven by the similarity of tasks, while in our case early- and late-game states are far from similar, which is why we believe inverse scaling is caused by indirect interaction between tasks through forgetting.
@@
The second limitation is the use of game states as proxies for independent task quanta. Oware inverse scaling cannot be solved simply by removing all late turns from the training distribution, possibly because the group of tasks that models overfit on only partially correlates with accuracy on late-game states. Fig. 7 misses a large chunk of information about the dataset due to the large number of states that only appear once in training. These states, at the tail-end of the frequency curve, have low individual frequency but are the majority of states encountered in training. In appendix F we visualize data frequency over a different axis and show that the tail-end of the state distribution is crucial for accurately calculating frequency scaling along other axes. A better grouping of task quanta might allow one to know which states to exclude from training in order to prevent inverse scaling.
@@ 8 Discussion
While this work touches several aspects of scaling law theory, we still lack a full model of RL scaling that can fully connect state distribution to Elo scaling laws. We provided clear evidence to the effect of Zipf’s law on loss, but loss is only a proxy for performance and the multi-agent Elo metric is connected to it in a non-trivial way. A significant problem with applying the quantization model to explain scaling laws is the difficulty of identifying task quanta, both for LLMs and RL, which is why we use game states as a proxy of tasks. Identifying the tasks or concepts that AlphaZero learns is a challenging interpretability problem, mostly limited to human game knowledge (McGrath et al., 2022; Schut et al., 2023).
We hope the analysis presented here will help advance the understanding of RL scaling. RL methods are often challenging to scale, and understanding why is a prominent topic in the field today (Kumar et al., 2020; Lyle et al., 2022; Sokar et al., 2023; Farebrother et al., 2024). Our results hint at a possibility for improved RL algorithms using a curriculum informed by the frequency distribution, similar to recent attempts in supervised learning (Ruoss et al., 2024).
Acknowledgments
We thank Eric J. Michaud for insightful feedback and discussions.
Appendix A Background: AlphaZero
AlphaZero is an RL algorithm for learning two-player zero-sum games, learning solely from self play (Silver et al., 2017a). The algorithm performs a Monte-Carlo tree search on the tree of future moves, guided by a neural network. The neural net receives the current board state as input, providing an estimate of its value, $v$, together with a policy prior, $\bf{p}$. The loss function is composed of separate elements for $v$ and $\bf{p}$:
@@ Appendix A Background: AlphaZero
During training, the agent visits board states $s$, having ground-truth labels $z$, for which it generates estimates $v$ and $\bm{\pi}$ for the state value and target prior, respectively. The MSE value-loss term $\left(z-v\right)^{2}$ pushes $v$ to fit the observed game outcome $z\in\{1,0,-1\}$, representing win, draw and loss respectively. The cross-entropy policy loss $-\bm{\pi}^{\top}\log\bf{p}$ pushes $\bf{p}$ to fit the actual policy $\bm{\pi}$ generated by the tree search. AlphaZero is known to follow scaling laws in certain games (Neumann & Gros, 2022; Jones, 2021).
Appendix B Zipf’s law variation
B.1 Zipf-exponent variation among agents
@@
In Fig. 9 we plot the state-frequency distributions of different games, plotting the combined distribution of many agents in the left-most column. We do this by counting the total frequency of states across games played by several different-sized agents. Each environment has its own Zipf exponent, but all exponents lie in the range $[0.75,1.05]$. For Oware, we fit the part of the curve that comes after the bump caused by late-game states, see appendix E. Checkers also exhibits a bump in the distribution, albeit harder to see since it starts at a much higher rank.
@@
Another set of Zipf curves can be obtained by counting state frequencies only in games played by a single agent. In each column after the left-most one, we plot the distribution generated by an agent with a different number of neural-net parameters. The standard deviation of the exponent among agents is low in all environments, standing at $0.01$ or lower for Connect Four, Pentago and Checkers, and at $0.02$ for Oware. The uniformity of the exponent across different agents agrees with prior observations on human games: the opening frequencies of human Go games also follow Zipf’s law, with the same exponent fitting both amateur games and games played at prestigious Go tournaments (Georgeot & Giraud, 2012).
B.2 Zipf’s law in non-uniformly-random games
Zipf’s law in games arises from their branching-tree structure. In section 5.1 we discuss the origin of Zipf’s law, showing that random toy-model games create a frequency distribution of plateaus centered on a power-law with exponent $\alpha=1$. We show that this distribution can be smoothed into a Zipf’s law by adding complicated game rules, but keeping the games random.

### v2

@@ 7 Analyzing how loss scales with Zipf’s law
In agreement with the quantization model michaud2023quantization , we find that AlphaZero models reduce their loss on game states in descending order of frequency. In Fig. 4 we plot the value-head loss on Connect Four training data in the limit of abundant compute, tracking the loss of each agent on its self-play games. [^3] The largest nets approach the optimal play limit neumann2022scaling , causing loss to stagnate as model size increases to the largest values.
Figure 4: Connect Four value-loss [^4] Note that AlphaZero’s value head is not equivalent to the value net of actor-critic methods, see appendix F.scaling with rank. A: Average loss on each agent’s own training data. Loss steadily increases with rank. Larger agents achieve better loss. B: Loss on a dataset annotated by a game solver. Larger models are closer to the ground truth values of optimal play. The overall trend of loss increasing with rank is surprising, when one considers that the complexity of states decreases with rank. See appendix E for detail on error bars. C: The time needed to evaluate the same states using alpha-beta pruning connect4solver drops on an exponential scale with rank (mean and standard deviation are geometric).
Comparison to the quantization model
The quantization hypothesis assumes models learn tasks in a binary way, sharply reducing the loss on each task in descending order of frequency. As a function of neural net size $N$, this assumed loss function is a step function that changes from low to high loss at rank $n\propto N$ michaud2023quantization . In comparison, we find that AlphaZero loss increases smoothly with rank rather than resembling a step function, likely because board positions are not independent task quanta.
In Fig. 4B we plot the value loss $L_{\rm value}=\left(z-v\right)^{2}$ on ground-truth labels $z$, calculated by an alpha-beta pruning solver connect4solver . We see that larger agents have significantly better ground-truth loss over smaller agents, confirming that increasing model size improves the absolute value loss. Ground-truth value loss is directly related to better performance, similar to test loss in LLMs that is known to correlate with accuracy on downstream tasks brown2020language ; srivastava2022beyond ; wei2022emergent .
@@
It is surprising that absolute value loss (Fig. 4B) tends to increase with state rank, since higher rank states are easier to model, meaning loss increases as state complexity decreases. Low-loss states are the hardest to model using conventional methods, as we show in Fig. 4C for alpha-beta pruning knuth1975analysis . The most frequent states have a larger state tree branching out from them, making it harder for search-based algorithms like MCTS to tackle them. As a result, the CPU time required to calculate the ground truth value of a state decreases steadily with rank on an exponential scale.
Figure 5: Inverse scaling in AlphaZero. A: Agents playing Checkers and Oware follow a size scaling law that abruptly changes direction, when large models fail to utilize their capacity. The scaling curve does not flip due to approaching the perfect-play ceiling: training a suite of Oware agents with different hyperparameters (light green), they extend to higher Elo scores. Error bars are one standard deviation. B: Oware games played by AlphaZero follow Zipf’s law, interrupted by a small plateau. This is caused by high-frequency late-game states, present due to the unusual tree-structure of Oware, see Fig. 6. C: Checkers shares the same tree structure, but the effect on the Zipf curve is less visible.
@@ 8 Connecting inverse scaling to the game structure
Inverse scaling, i.e. performance scaling negatively with training resources, is known to occur in LLM training mckenzie2023inverse . RL models can similarly fail to scale reliably, sometimes decreasing in performance as neural-network size and compute budget are increased. We see a clear example of inverse scaling for AlphaZero agents trained on Checkers and Oware, as shown in Fig. 5A. These agents follow a steady scaling curve that abruptly breaks down beyond a certain neural network size, where Elo starts to scale negatively with size.
@@ 8.1 Inverse scaling coincides with a frequency-distribution anomaly
Figure 6: Turn distribution anomaly coincides with inverse scaling. A: When the game tree spreads exponentially, e.g. in Connect Four and Pentago, turn number correlates with state frequency. However, in games with inverse scaling (Oware and Checkers) the game tree converges to a comparatively small set of endings, resulting in high-frequency late-game states. B: The fraction of late-game states (above dashed line in Oware and Checkers) rises either sharply or gradually with rank.
@@
States above and below the dashed line in Fig. 6A are clearly distinguishable when visualized, see examples in appendix G. Late-game states above the dashed line are positions close to the end of a game, when the board is mostly empty, while early-game states below the line are opening moves, i.e. small variations of the initial board state.
@@
The prevalence of high-frequency late-game states in Oware and Checkers data is a direct result of the game rules, which shape the game-tree structure. In both games, pieces are successively removed from the board, ending the game when one of the players has no pieces left (infinite-loop draws are rare). This win condition produces game trees that expand and then contract: starting from a predetermined position, expanding exponentially every move, but then contracting again towards the end of the game as the number of pieces decreases and with it the number of possible board-state permutations. The probability that a game ends in a frequently-visited board state is high.
@@ 8.2 Relation to inverse scaling
As we show in Fig. 6, late-game states (above the dashed line) dominate the Oware distribution at higher ranks, in a sharp transition from $0\%$ to $80\%$ of all states. This transition is drastic enough to visibly skew the frequency distribution seen in Fig. 5B, bending the Zipf power law. The distribution shift causes Oware value loss to spike, as we show in Fig. 7A. The loss of all but the largest models jumps by up to an order of magnitude at the transition point, showing that models that fit early-game states well have a hard time fitting late-game states. This counter-intuitive behavior of getting worse loss on easier states, which we discussed in section 7, is even more surprising here: Oware states after the transition point are predominantly end-game configurations, and should be significantly easier to model than early-game states using search algorithms.
The quantization model provides an explanation for the sharp rise in loss. As models fit the most frequent states, smaller models do not have enough capacity to additionally model less frequent states accurately. The large qualitative difference between early- and late-game states makes the difference in loss especially pronounced. Interestingly, larger agents converge on a smooth loss curve similar to that of Connect Four, at the cost of higher loss on early-game states.
@@ 8.3 Why do large models overfit?
The degradation of large model performance on early-game states is an example where RL scaling diverges from the quantization model. The quantization model explains how larger models could fit more tasks, without forgetting higher-frequency tasks. In RL, we see that fitting less-frequent states coincides with lower performance on the most frequent states, when states split into two distinct groups. The difference between supervised learning and RL scaling might be caused by a shift of the target distribution during training, which is known to cause training failure, e.g. through dormant neurons sokar2023dormant .
Figure 8: Large models forget knowledge on early-game states. A: The loss of the largest Oware agent improves (on average) during training. B: On early-game states, loss initially goes down, then stops improving and eventually worsens towards the end of training. C: Loss on late-game states steadily improves during training. The fact that large nets fit early-game states but then forget them suggests that target non-stationarity might be the cause of inverse scaling.
@@
A possible explanation of this selective forgetting is a selective shift in the distribution of the target labels, i.e. the value of each state. Sokar et al. (2023) sokar2023dormant have shown that changes to the target distribution in RL can harm performance by creating dormant neurons. Unlike early-game states, late-game states have relatively fixed label distributions since the outcome is already decided by that point in the game. A hypothetical mechanism of inverse scaling then follows: 1) Models fit states by order of frequency, gradually expanding the fraction of late-game states in the pool of learned states. 2) As new states are optimized, higher-frequency states learned earlier can be forgotten. 3) In Oware, late-game states are immune to forgetting due to their strongly fixed labels. 4) Large Oware models are trapped in local optima, fitting many end-game states well but forgetting a smaller fraction of crucial early-game states. See appendix H for a more in-depth explanation.
@@
We note a certain similarity between this inverse scaling phenomenon and one of the causes for inverse scaling identified by McKenzie et al. (2023) mckenzie2023inverse , namely distractor tasks. Some LLM benchmarks require mastering a set of tasks $T$, but as model scale increases, models first learn a set of easier tasks $D$ that harm performance, leading to inverse scaling. Another similar concept is competing tasks, which can cause inverse scaling in a supervised learning setting ildiz2024understanding . While the concept of competing/distractor tasks bears resemblance to the Oware case, the mechanism behind these phenomena is different. These examples of inverse scaling in supervised learning are driven by the similarity of tasks, while in our case early- and late-game states are far from similar, which is why we believe inverse scaling is caused by indirect interaction between tasks through forgetting.
@@
The second limitation is the use of game states as proxies for independent task quanta. Oware inverse scaling cannot be solved simply by removing all late turns from the training distribution, possibly because the group of tasks that models tend to overfit only partially correlates with late-game states. Fig. 7 misses some of the dataset due to states that only appear once in training. These states are at the tail-end of the frequency curve, given their low individual frequencies, but they constitute the majority of states encountered in training. In appendix I we visualize data frequency over a different axis and show that the tail-end of the state distribution, which we removed for our above analysis, is crucial for accurately calculating frequency scaling along other axes. A better grouping of task quanta might allow one to know which states to exclude from training in order to prevent inverse scaling.
@@ 9 Discussion
While this work touches several aspects of scaling law theory, we still lack a full model of RL scaling that can fully connect state distribution to Elo scaling laws. We provided clear evidence of the effect of Zipf’s law on loss, but loss is only a proxy for performance and the multi-agent Elo metric is connected to it in a non-trivial way. A significant problem with applying the quantization model to explain scaling laws is the difficulty of identifying task quanta, both for LLMs and RL, which is why we use game states as a proxy of tasks. Identifying the tasks or concepts that AlphaZero learns is a challenging interpretability problem, mostly limited to human game knowledge mcgrath2022acquisition ; schut2023bridging .
We hope the analysis presented here will help advance the understanding of RL scaling. RL methods are often challenging to scale, and understanding why is a prominent topic in the field today kumar2020implicit ; lyle2022understanding ; sokar2023dormant ; farebrother2024stop . Our results hint at a possibility for improved RL algorithms using a curriculum informed by the frequency distribution, similar to recent attempts in supervised learning ruoss2024grandmaster .
Acknowledgments and Disclosure of Funding
We thank Eric J. Michaud and David J. Wu for insightful feedback and discussions.
Appendix A Additional Background
AlphaZero
AlphaZero is an RL algorithm for learning two-player zero-sum games, learning solely from self play [20]. The algorithm performs a Monte-Carlo tree search on the tree of future moves, guided by a neural network. The neural net receives the current board state as input, providing an estimate of its value, $v$, together with a policy prior, $\bf{p}$. The loss function is composed of separate elements for $v$ and $\bf{p}$:
@@ Appendix A Additional Background
During training, the agent visits board states $s$, having ground-truth labels $z$, for which it generates estimates $v$ and $\bm{\pi}$ for the state value and target prior, respectively. The MSE value-loss term $\left(z-v\right)^{2}$ pushes $v$ to fit the observed game outcome $z\in\{1,0,-1\}$, representing win, draw and loss respectively. The cross-entropy policy loss $\bm{\pi}^{\top}\log\bf{p}$ pushes $\bf{p}$ to fit the actual policy $\bm{\pi}$ generated by the tree search.
AlphaZero scaling laws
AlphaZero scaling is one of the earliest examples of neural power-law scaling in RL [21, 5]. Jones [21] showed that AlphaZero agents trained to play Hex improve their Elo score $r$ with the log of training compute $C$, $r\propto\log(C)$, with a coefficient that remains constant across board sizes. As they point out, this predictable scaling behavior allows one to plan the compute budget ahead of training to fit a desired performance level. Neumann & Gros [5] found similar compute scaling in agents trained on Connect Four and Pentago, as well as scaling with neural network size, and pointed out that these scaling curves are power laws of the Bradley-Terry playing strength $\gamma$ [35], since $r\propto\log_{10}(\gamma)$. Similar to LLM scaling laws, these scaling laws of AlphaZero allow one to know in advance the optimal model size that would achieve maximal Elo score, given the available training compute budget. An interesting result from Jones shows that training compute can also be traded off predictably for inference compute [21].
Appendix B Board game descriptions
We describe here the rules of all four games examined in this paper. All games are two-player, zero-sum, open information games, where players alternate turns between them moving pieces on a board.
Connect Four
The players alternate placing a disk of their color into one of seven vertical columns on a 6x7 grid, and the disk falls to the lowest unoccupied position in that column. A player wins by arranging four of their disks in a row, either vertically, horizontally, or diagonally. If all cells are filled without such an alignment, the game is drawn.
Pentago
Players alternate placing one marble of their color on any empty cell of a 6x6 board, which is divided into four 3x3 rotating quadrants. After placing, the player must rotate one quadrant by 90° either clockwise or counter-clockwise. A player wins immediately upon obtaining five marbles in a row (horizontal, vertical, or diagonal) either before or after rotation. If the board becomes full without five in a row, the game is drawn.
Oware
This game is played on a board with two rows of six houses and optionally store pits. Each of the twelve small houses starts with four seeds. Each player controls the six houses on their side. On their turn, a player selects one of their houses with seeds, removes all seeds, and sows them counter-clockwise, placing one seed per subsequent house (excluding stores and skipping the original house when it’s revisited). If the last seed lands in an opponent’s house and brings its count to exactly two or three, those seeds are captured into the player’s store. Capture continues backward from that house as long as the immediately preceding opponent’s houses also contain exactly two or three seeds. Capturing all of an opponent’s seeds is forbidden; in that case no capture is made. The game ends when a player captures 25 or more seeds, or both players capture 24 each (draw), or if no legal move exists; remaining seeds are then collected by their owners. In our analyses, a draw is also declared if the game lasts 1000 turns.
Checkers
The players each control 12 pieces on opposite dark squares of an 8x8 chess board. Normal pieces move diagonally forward one square to an unoccupied dark square. If an opponent’s piece is diagonally adjacent and the following square is empty, the adjacent piece may be jumped and removed; multiple jumps are allowed in sequence if available. When a piece reaches the farthest row on the opponent’s side, it is crowned as a king and gains the ability to move and jump both forward and backward. Play continues until one player has no legal move (draw) or a player lost all their pieces. We also draw the game if it lasts too long.
Appendix C Zipf’s law variation
C.1 Zipf-exponent variation among agents
@@
In Fig. 9 we plot the state-frequency distributions of different games, plotting the combined distribution of many agents in the left-most column. We do this by counting the total frequency of states across games played by several different-sized agents. Each environment has its own Zipf exponent, but all exponents lie in the range $[0.75,1.05]$. For Oware, we fit the part of the curve that comes after the bump caused by late-game states, see appendix G. Checkers also exhibits a bump in the distribution, albeit harder to see since it starts at a much higher rank.
@@
Another set of Zipf curves can be obtained by counting state frequencies only in games played by a single agent. In each column after the left-most one, we plot the distribution generated by an agent with a different number of neural-net parameters. The standard deviation of the exponent among agents is low in all environments, standing at $0.01$ or lower for Connect Four, Pentago and Checkers, and at $0.02$ for Oware. The uniformity of the exponent across different agents agrees with prior observations on human games: the opening frequencies of human Go games also follow Zipf’s law, with the same exponent fitting both amateur games and games played at prestigious Go tournaments [30].
C.2 Zipf’s law in non-uniformly-random games
Zipf’s law in games arises from their branching-tree structure. In section 5 we discuss the origin of Zipf’s law, showing that random toy-model games create a frequency distribution of plateaus centered on a power-law with exponent $\alpha=1$. We show that this distribution can be smoothed into a Zipf’s law by adding complicated game rules, but keeping the games random.